OpenStudy (anonymous):

Write the equation of the line that passes through (2, 5) with a slope of 3 in point-slope form

Directrix (directrix):

y - 5 = 3(x - 2)

Directrix (directrix):

y = 3( x - 2) + 5 Line (point - slope form) (Coordinate Geometry) Definition: A straight line on the coordinate plane can be described by the equation y = m(x-Px) + Py where m is the slope of the line and Px, Py are the coordinates of a given point on the line .

Directrix (directrix):

http://www.mathopenref.com/coordequationps.html